# SonicSieve: Bringing Directional Speech Extraction to Smartphones Using Acoustic Microstructures

### v2

@@ ABSTRACT.
conference: ACM Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spaincopyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXX
@@ 1. INTRODUCTION
We build our end-to-end system on a smartphone with a low-cost wired earphone and a 3D-printed microstructure. Our results are as follows:
@@ 2. RELATED WORK
Acoustic sensing using microstructures. Acoustic microstructures have been designed for a variety of applications, including spatial sensing (Garg et al., 2021; Bai et al., 2022), sound absorption (Arjunan et al., 2024; Dupont et al., 2018), acoustic filtering (Li et al., 2016), and biometric identification (Yang et al., 2023). Owlet (Garg et al., 2021) designed a 3D-printed microstructure that embedded directional cues into incoming sounds, enabling direction-of-arrival (DoA) estimation and sound localization. SPiDR (Bai et al., 2022) extended this design by using microstructures to project and capture spatially coded signals to generate depth maps of nearby objects. EarCase (Li et al., 2023) adapted microstructures for smartphones, embedding them into a custom case for DoA estimation. However, its design required a new custom case for each device, limiting its ability to scale across multiple phone models. Our work differs in two key ways. First, our microstructure is designed for low-cost wired earphones, and is compatible with different smartphones. Second, our work is focused on directional speech extraction by leveraging spatial cues to selectively amplify and suppress speech based on their direction. This enables a broader set of spatial audio interactions beyond what is achievable through DoA alone.
@@
AI-enabled spatial sensing systems. These systems have focused on speech enhancement (Pascual et al., 2017; Chatterjee et al., 2022), sound localization (Shimada et al., 2021; Nguyen et al., 2022), subject-aware sound classification (Zhang et al., 2024; Wang et al., 2024b; Zhang et al., 2023) and speech separation (Jenrungrot et al., 2020; Wang et al., 2023). Recent works have leveraged spatial information to enable new capabilities on smart devices. ClearBuds (Chatterjee et al., 2022) leverages the placement of wireless earbuds to enable real-time speech enhancement for telephony applications. Cone of Silence (Jenrungrot et al., 2020) extracts speech from multiple target speakers at different locations. Sound Bubble (Chen et al., 2024) isolates audio within a defined bubble radius, suppressing sound from outside. ReZero (Gu and Luo, 2024) extracts all sounds within a user-defined spatial region—angular, spherical, or conical. However, these systems require either a fixed-wearable device like earbuds (Cao et al., 2020), large microphone arrays with at least six elements, or are not designed for causal, real-time use. In contrast, our work targets directional speech extraction in ad hoc scenarios such as meetings, where the positions of target speakers are unknown in advance. Additionally, we focus on smartphones, which, unlike microphone arrays, are mobile devices that people carry throughout the day.
Spatial sensing on smartphones. Prior work has focused on estimating DoA (Küçük et al., 2019; Zhou et al., 2024; Tokgöz et al., 2020), and beamforming (Gaubitch et al., 2014) on smartphones using statistical algorithms and neural networks. However, the performance of these systems is fundamentally limited by the number of microphones. Ad hoc microphone arrays composed of smartphones have also been proposed for spatial sensing (Sur et al., 2014). These systems emulate larger arrays by synchronizing multiple devices. However, they rely on complex synchronization protocols such as aligning smartphone CPU clocks using WiFi beacon timestamps, synchronizing audio I/O clocks through a backend server, or using time-of-arrival techniques (Li et al., 2017a; Peng et al., 2007), which can make deployment in real-world scenarios challenging.
@@ 3. SONIC SIEVE
This section outlines the technical approach of SonicSieve. We start with a high-level overview of how our system uses an acoustic microstructure for directional speech extraction. Then, we will detail the techniques for optimizing this microstructure, integrating it with a smartphone, using a neural network to extract speech in real-time, and real-world dataset collections.
@@
This figure shows two frequency response graphs labeled ”(a) Without Microstructure” and ”(b) With Microstructure” displaying amplitude of frequency response $M_{\theta}(f)$ (y-axis, ranging from -25 to 25) versus frequency in Hz (x-axis, ranging from 1000 to 8000 Hz). In graph (a) Without Microstructure, four colored lines representing different direction of arrival (0°, 60°, 120°, and 180°) show nearly flat, overlapping frequency responses with minimal variation between angles, all hovering around 0 amplitude across the frequency spectrum. In contrast, graph (b) With Microstructure shows the same four angles producing dramatically different frequency response patterns. The lines show significant peaks and valleys, particularly around 3000 and 6000 Hz where the greatest variation occurs. The 0° line (orange) and 60° line (red) show pronounced peaks reaching amplitudes of approximately 15 and 10 respectively, while the 120° line (green) shows a deep trough reaching approximately -15 amplitude in the same frequency region. The 180° line (blue) exhibits multiple smaller peaks and valleys throughout the spectrum.
[…]
@@ 3.1. System Overview
Traditional directional speech extraction relies on microphone arrays that use beamforming to isolate sound from specific directions. These systems work by analyzing timing and amplitude differences across multiple microphones—when sound arrives from a particular direction $\theta$, each microphone in the array receives the signal at slightly different times and intensities due to their spatial arrangement. The system can be modeled as:
@@
where $\mathbf{a}(f,\theta)$ encodes the directional information based on known geometric delays between microphone channels, and $\mathbf{u}(f,t)$ represents interference and noise. Beamforming algorithms then apply spatial filtering to enhance signals from the target direction while suppressing others.
@@
Instead of encoding the spatial information through microphone array geometry, SonicSieve leverages a passive acoustic microstructure that directly embeds direction-dependent spatial cues into the signal. Such a design is inspired by nature’s directional hearing system. For example, humans can localize and separated sound effectively with just two ears, which is due to the geometry of the head and ears creating distinctive reflecting and scattering effects that vary with sound angle (Van Opstal, 2016; Xie, 2013). Similarly, owls achieve exceptional sound localization through asymmetric ear structures (Coles and Guppy, 1988).
Specifically, our acoustic microstructure applies a direction-specific filter $M_{\theta}(f)$ to a sound source arriving at the microstructure from a relative angle $\theta$ and propagating through the internal structure. We present an example in Fig. 2(b) of the direction-specific filter $M_{\theta}(f)$ at four different angles without and with the microstructure. The plot shows that the microstructure can introduce substantially greater variation across different directions of arrival.
@@
However, with only single-channel audio, we can not isolate the spatial filter $M_{\theta}(f)$ from the signal source $s(f,t)$. Thus, we incorporate a secondary reference microphone placed outside the microstructure. The reference microphone records the sound unmodified by the microstructure simultaneously:
@@
We note that the received signal at the reference microphone from the target source $s^{\prime}(f,t)$ differs slightly from the one at the inside microphone $s(f,t)$, since the two microphones are not positioned at the exact same location, such that the arriving signals are propagated through slightly different environmental reflections. To enable robust estimation of the spatial cues across different environments, the two microphones need to be positioned as close as possible.
@@ 3.3. Integration of acoustic microstructure to smartphones
Figure 8. Effect of separation between earphone’s in-line mic and phone’s top mic on spectral variation. We compute the ratio of the received signals at the two microphones ($s_{ratio}(f)$) and estimate the variation across environments (lower is better). When the mics are closely spaced, there is less spectral variation.
[…]
@@ 3.4. Directional speech extraction network
Our network design adapts a state-of-the-art real-time sound extraction architecture (Chen et al., 2024) for angular-based directional speech extraction. The system processes short audio chunks (8 ms) to enable real-time applications such as live audio streaming and meeting transcription on smartphones.
@@ 3.5. Real-world dataset
Training Methodology. We process audio in 8 ms chunks (192 samples at 24 kHz) with 4 ms lookahead for real-time performance. The STFT uses a window size of 288 samples with 192-sample shifts, and we apply 96-sample zero padding at signal boundaries.
We optimize using a Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) loss from the Asteroid package (Pariente et al., 2020):
@@
We train for 150 epochs using the Adam optimizer (Kingma and Ba, 2014). The learning rate starts at $5\texttimes 10^{-4}$, increases to $5\texttimes 10^{-3}$ over 10 epochs, remains constant for 20 epochs, then decays by 0.95 every 2 epochs. We apply data augmentation (Time Shift, Time Stretch, Frequency Mask, and Gain Perturbation) with 30% probability for each method (Wei et al., 2020).
@@ 4.1. Model results
Figure 11. System performance compared to the baselines without microstructure (a) Effect of selected sector index on SI-SDR improvement (b) SI-SDR improvement as a function of the number of simultaneously selected sectors.
[…]
This figure presents three graphs evaluating system performance through SI-SDR (Scale-Invariant Signal-to-Distortion Ratio) measurements across different conditions: (a) A line graph showing the effect of selected sector index (x-axis, numbered 1-6) on SI-SDR improvement in dB (y-axis, ranging 2-7 dB). Two lines compare performance: With microstructure (light green line with circles): Shows significantly higher SI-SDR values (between 3.6-6.8 dB) with a U-shaped pattern where sectors 1 and 6 perform best. The SI-SDR number for each sector is 6.4, 4.7, 4.4, 4.1, 3.6, 6.8. Without microstructure (dark blue line with squares): Shows consistently lower performance (between 1.9-2.8 dB) with a slight downward trend as sector index increases. The SI-SDR number for each sector is 2.7, 2.0, 2.5, 2.4, 2.1, 1.9. (b) A bar graph comparing SI-SDR improvement (y-axis, in dB) across different numbers of simultaneously queried sectors (x-axis, 1-3 sectors). For each number of sectors, two bars compare: With microstructure (green bars): Shows 5.0 dB for 1 sector, 4.2 dB for 2 sectors, and 3.9 dB for 3 sectors. Without microstructure (blue bars): Shows substantially lower values of 2.3 dB, 1.9 dB, and 1.7 dB respectively.
@@
Fig. 12(a) shows the relationship between the input SI-SDR of the speech samples in our dataset and the output SI-SDR after processing with our model when 1, 2, or 3 sectors are selected. The diagonal dashed line represents zero improvement, where the output SI-SDR equals the input. Points above this line indicate enhanced speech quality after processing. As evident from the figure, the vast majority of our test samples (86.6%) demonstrate positive enhancement, with output SI-SDR values consistently higher than their original values. Notably, the one-sector selection configuration (represented by red dots) shows the fewest samples (9.9%) falling below the zero-improvement line compared to two- and three-sector selections (15.2% and 16.3%, respectively), confirming our earlier observation that the task becomes more challenging as the number of selected sectors increases.
Figure 12. Overall system performance. (a) Input vs. output SI-SDR of our model across all samples, grouped by 1-sector, 2-sector, and 3-sector selections. Points above the dashed line indicate cases where SI-SDR improved. (b) Comparison of SI-SDR improvement across microphone arrays of varying sizes versus our SonicSieve system.
[…]
@@
(a) A scatter plot showing input vs. output SI-SDR (x-axis: Original SI-SDR ranging from -25 to 15 dB; y-axis: Output SI-SDR ranging from -30 to 20 dB). Points are color-coded by query type: 1 Sector Query (red dots), 2 Sectors Query (blue dots), 3 Sectors Query (orange dots). A diagonal dashed line indicates ”Zero improvement” reference. Most data points cluster above this line (especially in the -5 to 5 dB original SI-SDR range), demonstrating positive enhancement, with the upper-right quadrant labeled ”Enhancement” in green. (b) A bar plot compares SI-SDR improvement (y-axis, in dB ranging from 2-5.2 dB) across microphone arrays of varying sizes versus the SonicSieve system. The x-axis shows different configurations: 2-mic:  2.0 dB, 3-mic:  3.7 dB, 4-mic:  4.0 dB, 5-mic:  4.1 dB, 2-mic (SonicSieve):  4.5 dB (shown with green diagonal stripes pattern), 6-mic:  5.1 dB. This demonstrates that the 2-mic SonicSieve system achieves performance comparable to arrays with more microphones, outperforming the 5-mic arrays while using fewer hardware resources.
@@
Comparison against microphone array. We compared the end-to-end performance of SonicSieve against a conventional microphone array (UMA-8 USB mic array V2.0) as a baseline. For this baseline, we leveraged 2 to 6 channels from the array and trained separate neural networks for each configuration. All baseline models were trained on the same dataset with the same architecture as SonicSieve, except for minor modifications to accommodate the varying number of input channels.
@@
Fig. 12(b) shows the SI-SDRi, computed as the difference in SI-SDR between the sound mixture at the reference microphone and the enhanced speech output from the model. As the number of microphone channels increases from 2 to 6, SI-SDRi improves from 2.1 dB to 5.1 dB, due to the increased spatial diversity. SonicSieve achieves an SI-SDRi of 4.4 dB, outperforming the 5-microphone configuration (4.1 dB), and approaching the performance of the 6-microphone configuration (5.1 dB), demonstrating the effectiveness of our whole system.
@@
[…]

### v3

@@ ABSTRACT.
Directional speech extraction, Acoustic microstructures, Spatial sensing, Audio interfaces, Machine Learning
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790376isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Sound-based input / outputccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Computing methodologies Machine learning
@@ 1. INTRODUCTION
We build our end-to-end system on a smartphone with a low-cost wired earphone and a 3D-printed microstructure. This prototype therefore targets smartphones that can physically connect to a wired earphone and currently relies on an app-guided alignment procedure between the microstructure and the phone’s top microphone. We revisit these design trade-offs in our discussion section. Our results are as follows:
@@ 2. RELATED WORK
Acoustic sensing using microstructures. Acoustic microstructures have been designed for a variety of applications, including spatial sensing (Garg et al., 2021; Bai et al., 2022), sound absorption (Arjunan et al., 2024; Dupont et al., 2018), acoustic filtering (Li et al., 2016), and biometric identification (Yang et al., 2023). Owlet (Garg et al., 2021) designed a 3D-printed microstructure that embedded directional cues into incoming sounds, enabling direction-of-arrival (DoA) estimation and sound localization. SPiDR (Bai et al., 2022) extended this design by using microstructures to project and capture spatially coded signals to generate depth maps of nearby objects. EarCase (Li et al., 2023) adapted microstructures for smartphones, embedding them into a custom case for DoA estimation. However, its design required a new custom case for each device, limiting its ability to scale across multiple phone models and microphone layouts. In addition, case-based designs fix the relative placement between the microphones, which can make it harder to optimize for close microphone spacing that improves robustness across environments. Our work differs in two key ways. First, our microstructure is designed for low-cost wired earphones, and is compatible with different smartphones. This form factor allows the in-line microphone to be positioned close to the phone’s built-in microphone, which empirically improves reliability under changing room acoustics (Sec. 3.3). Second, our work is focused on directional speech extraction by leveraging spatial cues to selectively amplify and suppress speech based on their direction. This enables a broader set of spatial audio interactions beyond what is achievable through DoA alone.
@@
AI-enabled spatial-aware audio processing. These systems have focused on speech enhancement (Pascual et al., 2017; Chatterjee et al., 2022), sound localization (Shimada et al., 2021; Nguyen et al., 2022; Yuan et al., 2025b), sound classification (Zhang et al., 2024; Wang et al., 2024b; Zhang et al., 2023; Yuan et al., 2024b) and speech separation (Jenrungrot et al., 2020; Wang et al., 2023). Recent works have leveraged spatial information to enable new capabilities on smart devices. Cone of Silence (Jenrungrot et al., 2020) extracts speech from multiple target speakers at different locations. ReZero (Gu and Luo, 2024) extracts all sounds within a user-defined spatial region—angular, spherical, or conical. However, most of these systems either require large microphone arrays with at least six elements, or are not designed for causal, real-time use. ClearBuds (Chatterjee et al., 2022) and Sound Bubble (Chen et al., 2024) achieve causal, real-time performance using wearable devices. ClearBuds leverages wireless earbuds (Cao et al., 2020) for real-time speech enhancement in telephony, while Sound Bubble isolates audio within a defined radius. However, these systems are designed to enhance the wearer’s hearing in human-human communication. In contrast, our work targets smartphone-based directional speech extraction for general audio recording and streaming applications, such as meeting transcription and voice assistants. Smartphones are well-suited for ad-hoc application scenarios as they are mobile devices that people carry throughout the day.
Spatial sensing on smartphones. Prior work has focused on estimating DoA (Küçük et al., 2019; Zhou et al., 2024; Tokgöz et al., 2020), and beamforming (Gaubitch et al., 2014) on smartphones using statistical algorithms and neural networks. However, the performance of these systems is fundamentally limited by the number of microphones. Ad hoc microphone arrays composed of smartphones have also been proposed for spatial sensing (Sur et al., 2014). These systems emulate larger arrays by synchronizing multiple devices. However, they rely on complex synchronization protocols such as aligning smartphone CPU clocks using WiFi beacon timestamps, synchronizing audio I/O clocks through a backend server, or using time-of-arrival techniques (Li et al., 2017b; Peng et al., 2007), which can make deployment in real-world scenarios challenging.
@@ 3. SONIC SIEVE
This section outlines the technical approach of SonicSieve. Our goal is to let users pick a direction on a phone interface and have the phone amplify that speaker in real time, using a low-cost add-on rather than specialized microphone arrays. We start with a high-level overview of how our system uses an acoustic microstructure for directional speech extraction. Then, we will detail the techniques for optimizing this microstructure, integrating it with a smartphone, using a neural network to extract speech in real-time, and real-world dataset collections.
@@ 3.1. System Overview
Traditional directional speech extraction relies on microphone arrays that use beamforming to isolate sound from specific directions. These arrays are common in dedicated devices (e.g., smart speakers), but are less accessible on everyday smartphones due to hardware constraints. These systems work by analyzing timing and amplitude differences across multiple microphones—when sound arrives from a particular direction $\theta$, each microphone in the array receives the signal at slightly different times and intensities due to their spatial arrangement. The system can be modeled as:
@@
where $\mathbf{a}(f,\theta)$ encodes the directional information based on known geometric delays between microphone channels, and $\mathbf{u}(f,t)$ represents interference and noise. Intuitively, $\mathbf{a}(f,\theta)$ describes how a sound from direction $\theta$ “shows up” across microphones, enabling spatial filtering. Beamforming algorithms then apply spatial filtering to enhance signals from the target direction while suppressing others.
@@
Instead of encoding the spatial information through microphone array geometry, SonicSieve leverages a passive acoustic microstructure that directly embeds direction-dependent spatial cues into the signal. Such a design is inspired by nature’s directional hearing system. For example, humans can localize and separated sound effectively with just two ears, which is due to the geometry of the head and ears creating distinctive reflecting and scattering effects that vary with sound angle (Van Opstal, 2016; Xie, 2013). Similarly, owls achieve exceptional sound localization through asymmetric ear structures (Coles and Guppy, 1988). In practice, our microstructure acts like an acoustic lens placed over a microphone: it reshapes incoming sound differently depending on where it comes from (Fig. 3).
Specifically, our acoustic microstructure applies a direction-specific filter $M_{\theta}(f)$ to a sound source arriving at the microstructure from a relative angle $\theta$ and propagating through the internal structure. We present an example in Fig. 2 of the direction-specific filter $M_{\theta}(f)$ at four different angles without and with the microstructure. The microstructure introduces substantially greater variation across different directions of arrival—without it, the signals look almost identical across angles; with it, each angle leaves a more distinctive spectral signature that a model can learn to exploit.
@@
However, with only single-channel audio, we can not reliably isolate the spatial filter $M_{\theta}(f)$ in real-world settings. The challenge is that different sound sources have different intrinsic spectra, and room reflections further modify the signal. With one channel, the model cannot distinguish whether a spectral pattern comes from the microstructure’s directional filtering, the source itself, or the room. A single-mic system might work for fixed sources in controlled environments, but cannot generalize across diverse real-world conditions.
Thus, we incorporate a secondary reference microphone placed outside the microstructure. This reference gives the model a “before/after” view: one channel is shaped by the microstructure, and one is not. The reference microphone records the unmodified sound as:
@@
We note that the received signal at the reference microphone from the target source $s^{\prime}(f,t)$ differs slightly from the one at the inside microphone $s(f,t)$, since the two microphones are not positioned at the exact same location, such that the arriving signals are propagated through slightly different environmental reflections. To minimize this mismatch and enable robust estimation of spatial cues across different environments, we position the two microphones as close as practically feasible.
@@
In other words, given the two recordings and the user-selected sector(s), the network outputs only the speech coming from those directions.
@@ 3.3. Integration of acoustic microstructure to smartphones
Figure 8. Effect of separation between earphone’s in-line mic and phone’s top mic on spectral variation. We compute the ratio of the signals at the two microphones ($s_{ratio}(f)$) and estimate the variation across environments (lower is better). The variations are lower when the mics are closely spaced.
@@ 3.4. Directional speech extraction network
Our network design adapts a state-of-the-art real-time sound separation architecture TF-GridNet (Wang et al., 2023; Chen et al., 2024) for angular-based directional speech extraction. The system processes short audio chunks (8 ms) to enable real-time applications such as live audio streaming and meeting transcription on smartphones.
@@ 3.5. Real-world dataset
Training Methodology. We process audio in 8 ms chunks (192 samples at 24 kHz) with 4 ms lookahead for real-time performance. The STFT uses a window size of 288 samples with 192-sample shifts, and we apply 96-sample zero padding at signal boundaries. We optimize using a Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) loss from the Asteroid package (Pariente et al., 2020):
@@
We train for 150 epochs using the Adam optimizer (Kingma and Ba, 2014). The learning rate starts at $5\times 10^{-4}$, increases to $5\times 10^{-3}$ over 10 epochs, remains constant for 20 epochs, then decays by 0.95 every 2 epochs. We apply data augmentation (Time Shift, Time Stretch, Frequency Mask, and Gain Perturbation) with 30% probability for each method (Wei et al., 2020).
@@ 4.1. Model results
Figure 11. Overall System performance. (a) Effect of selected sector index on SI-SDR improvement. (b) SI-SDR improvement as a function of the number of simultaneously selected sectors. (c) Input vs. output SI-SDR of our model across all samples, grouped by 1-sector, 2-sector, and 3-sector selections. Points above the dashed line indicate cases where SI-SDR improved.
This figure presents three graphs evaluating system performance through SI-SDR (Scale-Invariant Signal-to-Distortion Ratio) measurements across different conditions: (a) A line graph showing the effect of selected sector index (x-axis, numbered 1-6) on SI-SDR improvement in dB (y-axis, ranging 2-7 dB). Two lines compare performance: With microstructure (light green line with circles): Shows significantly higher SI-SDR values (between 3.6-6.8 dB) with a U-shaped pattern where sectors 1 and 6 perform best. The SI-SDR number for each sector is 6.4, 4.7, 4.4, 4.1, 3.6, 6.8. Without microstructure (dark blue line with squares): Shows consistently lower performance (between 1.9-2.8 dB) with a slight downward trend as sector index increases. The SI-SDR number for each sector is 2.7, 2.0, 2.5, 2.4, 2.1, 1.9. (b) A bar graph comparing SI-SDR improvement (y-axis, in dB) across different numbers of simultaneously queried sectors (x-axis, 1-3 sectors). For each number of sectors, two bars compare: With microstructure (green bars): Shows 5.0 dB for 1 sector, 4.2 dB for 2 sectors, and 3.9 dB for 3 sectors. Without microstructure (blue bars): Shows substantially lower values of 2.3 dB, 1.9 dB, and 1.7 dB respectively. (c) A scatter plot showing input vs. output SI-SDR (x-axis: Original SI-SDR ranging from -25 to 15 dB; y-axis: Output SI-SDR ranging from -30 to 20 dB). Points are color-coded by query type: 1 Sector Query (red dots), 2 Sectors Query (blue dots), 3 Sectors Query (orange dots). A diagonal dashed line indicates ”Zero improvement” reference. Most data points cluster above this line (especially in the -5 to 5 dB original SI-SDR range), demonstrating positive enhancement, with the upper-right quadrant labeled ”Enhancement” in green.
@@
Fig. 11(c) shows the relationship between the input SI-SDR of the speech samples in our dataset and the output SI-SDR after processing with our model when 1, 2, or 3 sectors are selected. The diagonal dashed line represents zero improvement, where the output SI-SDR equals the input. Points above this line indicate enhanced speech quality after processing. As evident from the figure, the vast majority of our test samples (86.6%) demonstrate positive enhancement, with output SI-SDR values consistently higher than their original values. Notably, the one-sector selection configuration (represented by red dots) shows the fewest samples (9.9%) falling below the zero-improvement line compared to two- and three-sector selections (15.2% and 16.3%, respectively), confirming our earlier observation that the task becomes more challenging as the number of selected sectors increases.
Figure 12. Comparison of SI-SDR improvement across microphone arrays of varying sizes versus our SonicSieve system.
@@
A bar plot compares SI-SDR improvement (y-axis, in dB ranging from 2-5.2 dB) across microphone arrays of varying sizes versus the SonicSieve system. The x-axis shows different configurations: 2-mic:  2.0 dB, 3-mic:  3.7 dB, 4-mic:  4.0 dB, 5-mic:  4.1 dB, 2-mic (SonicSieve):  4.5 dB (shown with green diagonal stripes pattern), 6-mic:  5.1 dB. This demonstrates that the 2-mic SonicSieve system achieves performance comparable to arrays with more microphones, outperforming the 5-mic arrays while using fewer hardware resources.
@@
Comparison against microphone array and traditional beamforming. We compared the end-to-end performance of SonicSieve against a conventional microphone array (UMA-8 USB mic array V2.0) as a baseline. Our network design is adapted from TF-GridNet (Wang et al., 2023), a state-of-the-art real-time sound separation architecture that supports arbitrary numbers of input channels, enabling fair comparison across different microphone configurations. For this baseline, we used 2 to 6 channels from the array and trained separate neural networks for each configuration. All baseline models were trained on the same dataset with the same architecture as SonicSieve, except for minor modifications to accommodate the varying number of input channels.
@@
Fig.12 shows the SI-SDRi, computed as the difference in SI-SDR between the sound mixture at the reference microphone and the enhanced speech output from the model. As the number of microphone channels increases from 2 to 6, SI-SDRi improves from 2.1 dB to 5.1 dB, due to the increased spatial diversity. SonicSieve achieves an SI-SDRi of 4.4 dB, outperforming the 5-microphone configuration (4.1 dB), and approaching the performance of the 6-microphone configuration (5.1 dB). Notably, the 2-microphone neural baseline achieves only 2.1 dB SI-SDRi, demonstrating that SonicSieve’s 4.4 dB improvement stems from the microstructure’s ability to fundamentally enrich spatial cues beyond what microphone spacing alone provides.
We also evaluated the classical MVDR (Minimum Variance Distortionless Response) beamforming algorithm on the microphone array dataset. MVDR was configured to beamform toward the center of each selected sector (e.g., 15° for sector 1 spanning 0° to 30°). However, MVDR achieved less than 2 dB SI-SDRi even when using all 6 microphone channels. This poor performance stems from several fundamental limitations: (1) the coarse 30° sector resolution cannot provide the precise direction-of-arrival information that MVDR requires for effective spatial filtering, (2) MVDR’s linear formulation struggles to handle complex multipath propagation and reverberation, which diffuse the spatial cues, (3) MVDR cannot effectively separate speech from environmental noise when they originate from similar directions, as it relies solely on spatial cues rather than the richer acoustic signatures that neural models can exploit.
@@
Figure 14. User study measuring listening experience via mean opinion scores: (a) audibility of interfering speakers and background noise, (b) overall listening quality, (c) overall listening quality across the audio samples collected in different rooms
This figure displays results from a user study measuring listening experience through mean opinion scores on a 1-5 scale (higher is better), presented in three panels with error bars: (a) ”Noise Suppression Performance” compares three systems: 2-mic array with mean score of approximately 2.5 (error bars ranging from about 1.3 to 3.7), 5-mic array with mean score of approximately 3.5 (error bars ranging from about 2.1 to 4.9), and SonicSieve with mean score of approximately 3.3 (error bars ranging from about 2.1 to 4.4). (b) ”Overall Experience” compares the same three systems: 2-mic array with mean score of approximately 2.5 (error bars ranging from about 1.3 to 3.7), 5-mic array with mean score of approximately 3.3 (error bars ranging from about 2.1 to 4.5), and SonicSieve with mean score of approximately 3.4 (error bars ranging from about 2.1 to 4.7). (c) ”Overall Experience by Room” breaks down the overall listening quality across three different rooms (Room A, Room B, Room C), with grouped bars for each system: In Room A, 2-mic array scores approximately 2.2, 5-mic array approximately 3.3, and SonicSieve approximately $3.$ In Room B, 2-mic array scores approximately 2.4, 5-mic array approximately 3.3, and SonicSieve approximately $3.$ In Room C, 2-mic array scores approximately 2.5, 5-mic array approximately 3.5, and SonicSieve approximately 3.5. Panel (a) specifically measures audibility of interfering speakers and background noise (noise suppression capability), while panels (b) and (c) measure overall listening quality. The data shows that SonicSieve and the 5-mic array perform comparably across all metrics and consistently outperform the 2-mic array baseline in all rooms. The overlapping error bars indicate variability in user opinions, but the trends demonstrate clear preferences for the SonicSieve and 5-mic systems across different acoustic environments.